trying to find wing span of bird that is 8.75 lbs. with the equation of L=2.43*W^0.3326 in step by step method please explain
Answer by rfer(16322)

L=2.43*W^0.3326
L=2.43(8.75)^0.3326
L=2.48(2.0573675)
L=5
---------------------
first take 8.75 to the 0.3326 power
then multiply that answer by 2.48